OpenStudy (anonymous):

State whether the system of equations has one solution, no solutions, or infinite solutions a. y = x -3 -x + 3y = -9

OpenStudy (anonymous):

if you would draw a line for each of these functions, how many times do you think they would cross each other?

OpenStudy (anonymous):

I have absolutely no cue. Sorry.

OpenStudy (anonymous):

clue*

OpenStudy (anonymous):

@agent0smith he can explain this! with his smarticles:)

OpenStudy (anonymous):

Two straight lines will always cross once, unless they are parallel, than they will cross never.

OpenStudy (anonymous):

so all you have to do, is to see if they are parallel or not, if they are not, than they will cross, once.

OpenStudy (agent0smith):

You can either draw a graph, or try to solve the system to decide y = x -3 -x + 3y = -9 use substitution since you know y is equal to x-3, so replace the y in the second equation: -x + 3(x-3) = -9 see if you can solve that for x

OpenStudy (anonymous):

See i know smarticles when i see it:)